OpenStudy (superhelp101):

Solve x2 + 8x – 48 = 0 by completing the square.

OpenStudy (superhelp101):

@alg2help

OpenStudy (anonymous):

Ppsh, who uses completing the square

OpenStudy (anonymous):

well you can use the factor 12 and - 4 so you can get 8 so the simplified verion will be (x + 12)(x - 4)

OpenStudy (anonymous):

quadratic equation!!!!!!

OpenStudy (zzr0ck3r):

completing the square is by far the best imo for \(x^2+bx+c=0\) we have \((x+\frac{b}{2})^2=-c+(\frac{b}{2})^2\)

OpenStudy (anonymous):

No way, completing the squares is the most confusing way there is!

OpenStudy (anonymous):

first of all lets make the equation x^2 + 8x = 48 now we find (b/2)^2 which is 16 add 16 to both sides x^2 + 8x + 16 = 48 + 16 (x+4)^2 = 64 rad both sides: x+4 = +-8 x = 4 x = -12

OpenStudy (superhelp101):

lol thanks everyone! :)

OpenStudy (anonymous):

You're welcome - everyone

OpenStudy (superhelp101):

hehe :)

OpenStudy (zzr0ck3r):

if the middle term is even and there is no coefficient on the first term, and you practice a few times. You can solve these in your head.

OpenStudy (zzr0ck3r):

by completing the square....

OpenStudy (superhelp101):

oh ok

OpenStudy (zzr0ck3r):

:)

OpenStudy (superhelp101):

:)

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

:)